Problem – A silicon wafer is doped with 2×1016 boron and 1016 phosphorus atoms per cm3


.


Calculate n, p, and EF at room temperature assuming complete ionization. Repeat the same for 8


x 1015 boron atoms per cm3


and 1016 phosphorus atoms per cm3


.

1
Expert's answer
2022-02-01T15:33:48-0500

1)

"n_1=2\cdot 10^{16}," "p_1=5\cdot 10^3,"

"n_2=10^{16}," "p_2=10^4,"

2)

"n_1=8\cdot10^{15}," "p_1=1.25\cdot 10^4,"

"n_2=10^{16}," "p_2=10^4."